LABORATORY DETECTION AND MICROWAVE SPECTRUM of ScC_{2} RADICAL (X^{2}A_{1})

\underline{JIE MIN}, DEWAYNE T. HALFEN, LUCY. M. ZIURYS, Department of Chemistry and Biochemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721.

The pure rotational spectrum of ScC_{2} (X^{2}A_{1}) has been measured using Fourier transform microwave (FTMW) spectroscopy. This is the first study of ScC_{2} by any spectroscopic technique. In the FTMW system, the molecule was synthesized using the discharge assisted laser ablation source (DALAS) from a mixture of 0.25 % methane in argon and the ablation of a scandium rod. Spectra of the main isotopologue, ScC_{2}, as well as Sc^{13}C^{13}C and Sc^{13}C^{12}C, have been recorded in the frequency range of 4-60 GHz. Three rotational transitions have been measured, each exhibiting fine structure and hyperfine splittings due to scandium nuclear spin (I =7/2). The data have been analyzed with a ^{2}A_{1} Hamiltonian and rotational, spin-rotation constants and Sc and ^{13}C hyperfine parameters have been determined. The spectra are consistent with a triangular structure and a ^{2}A_{1} ground electronic state as predicted by theory. A precise structure will be presented.